OpenStudy (anonymous):

f(x) = x + (128) / (x^2+16). Identify all relative extrema. Here is my progress: After Deriving etc, I came up with: (x^2 + 16)^2 - 256x = 0 I tried to expand and got x^4 + 16x^4 + 256 - 256x = 0 Can someone help me solve this ? Thanks!

OpenStudy (abb0t):

\[f' = -\frac{ x^2+256x-16 }{ (x^2+16)^2 }\]

OpenStudy (abb0t):

To find the max, you set it equal to zero. Which should make the problem fairly easy to solve.

OpenStudy (anonymous):

Can you give me some assistance with this part, I always have a hard time :(

OpenStudy (abb0t):

HINT: . \[\frac{ ax }{ c } = 0 \] \[c \times 0 = 0\] hence \[ax =0\]

OpenStudy (anonymous):

Honestly I have no idea how to even solve it. :(. Can you solve it for me please?

OpenStudy (abb0t):

I already gave you the derivative. When you multiply the denominator by zero, you only get a quadratic! You don't know how to factor the quadratic? Is that where you're struggling?!

OpenStudy (abb0t):

Factor: \[-(x^2+256x-16) = 0\]

OpenStudy (anonymous):

Oh...I think I am starting to understand lol.

OpenStudy (abb0t):

\[(x \pm ?_1 ) (x \pm ?_2)\]

OpenStudy (anonymous):

So I should do what after " −(x2+256x−16)=0"

OpenStudy (abb0t):

factor it out! (x + ___ )( x + ___) solve for x.

OpenStudy (anonymous):

Uh..How do I factor that? my results dont make sense..

OpenStudy (anonymous):

what did u get >.>